# Microwave driven singlet-triplet qubits enabled by site-dependent g-tensors

### v1

@@ Abstract
Hole spin qubits are rapidly emerging as the workhorse of semiconducting quantum processors because of their large spin-orbit interaction, enabling fast all-electric operations at low power. However, spin-orbit interaction also causes non-uniformities in devices, resulting in locally varying qubit energies and site-dependent anisotropies. While these anisotropies can be used to drive single-spins, if not properly harnessed, they can hinder the path toward large-scale quantum processors. Here, we report on a qubit in planar germanium, where we control the anisotropy of two spins in a double quantum dot. By characterising microwave-driven singlet-triplet qubits, we show two distinct operating regimes depending on the magnetic field direction. For in-plane fields, the two spins are largely anisotropic, enabling to measure all the available transitions; coherence times exceeding 3 $\mu$s are extracted. For out-of-plane fields, they have an isotropic response but preserve the substantial energy difference required to address the singlet-triplet qubit. Even in this field direction, where the qubit lifetime is strongly affected by nuclear spins, we find 400 ns coherence times. Our work adds a valuable tool to investigate and harness the anisotropy of spin qubits and can be implemented in any large-scale NxN device, facilitating the path towards scalable quantum processors.
@@
Here, we make use of a double quantum dot device formed in a Ge/SiGe heterostructure to characterise the g-tensor and exchange anisotropy, key quantities for future scaling of semiconductor spin qubit quantum processors. We develop a simple protocol which allows to systematically characterize all the relevant spectroscopic features of two-hole systems at fixed magnetic field directions, including their relative quantization axes. For in-plane fields, we investigate microwave-driven singlet-triplet transitions –including the elusive $T_{-}$ to $T_{+}$ transition– and we show that such transitions are driven via the modulation of the exchange interaction and are just possible because of g-tensor dissimilarities. For out-of-plane fields, we demonstrate spin axes alignment and realize a $S-T_{0}$ microwave-driven qubit, which already for small fields has a large energy splitting and small transition matrix elements to the leakage states $T_{-}$ and $T_{+}$. This solves the problem of the non-orthogonal rotation axes, which have been limiting their implementation.
@@
We start by mapping out the amplitude of the signal at the Pauli Spin Blockade point as a function of the in-plane magnetic field $B_{y}$ and frequency of the microwave burst as seen in Fig. 1(c). We observe the appearance of three spin transitions, in contrast to the typically observed two lines  [25, 26, 27]. Those transitions are due to spin flips of charges localized in the two QDs. The faint line at the lowest frequency has half the frequency of the line indicated by the red arrow. This subharmonic transition is a consequence of a non-linear driving mechanism (see SFig. S9)  [28, 29].
@@
where $\theta=-\text{arctan2}(\sqrt{2}t_{c}/\varepsilon)/2$ is the mixing angle, $J=-\varepsilon/2+\sqrt{\varepsilon^{2}/4+8t_{c}^{2}}$ is the exchange interaction, with $t_{c}$ the tunnel coupling. $\bm{b}$ ($\delta\bm{b}$) is the addition (difference) of the Zeeman vectors $\mu_{B}\underline{g}_{i}\bm{B}/2$ of the two dots ($i=1,2$), where $\mu_{B}$ is the Bohr magneton and $\bm{B}$ is the magnetic field. The parallel and perpendicular projections of $\delta\bm{b}$ on $\bm{b}$ are denoted $\delta b_{\parallel}$ and $\delta b_{\perp}$, respectively (see Fig. 1(f)). For convenience, we introduce the corresponding dimensionless g-factors as $\bar{b}=\mu_{B}|(\underline{g}_{1}+\underline{g}_{2})\bm{B}|/2=\mu_{B}\bar{g}|\bm{B}|$, $\delta b_{\parallel}=\mu_{B}\delta g_{\parallel}|\bm{B}|$ and $\delta b_{\perp}=\mu_{B}\delta g_{\perp}|\bm{B}|$. It is important to note that $\bar{g}$, $\delta g_{\parallel}$ and $\delta g_{\perp}$ are effective g-factors, which also incorporate the spin-flip tunnelling effects  [34, 32, 33].
@@
We emphasize that an equivalent description appropriate for large B field comprises two aligned spins coupled by an anisotropic exchange interaction matrix $JR_{y}(-\delta\theta)$. The rotation angle is given by $\delta\theta=\text{arctan}\left[\delta g_{\perp}/(\bar{g}+\delta g_{\parallel})\right]+\text{arctan}\left[\delta g_{\perp}/(\bar{g}-\delta g_{\parallel})\right]$ (see Methods). From the above experimentally extracted values we find $\delta\theta=45^{\circ}$, highlighting the large degree of exchange anisotropy in our system. We note that large anisotropy is expected to be useful for resonant two qubit gates in Loss-DiVincenzo qubits and for leakage protected two-qubit gates in ST qubits.
@@
After the characterization of the spectroscopic features, we focus on extracting the mechanism which permits to drive these transitions. We study the dependence of the spin transitions as a function of detuning at a fixed magnetic field in Fig. 2(a). As the system is initialized in the ground state, the three observed lines correspond to the transitions indicated by the three arrows depicted in Fig. 1(e). From the $T_{-}-S$ transition, we can experimentally determine the position, $\varepsilon_{AC}=1.8$ meV, and size, $\Delta_{ST_{-}}=175$ MHz, of the $S-T_{-}$ anticrossing. As this point sets the turnover of the ground state, we identify two regimes. When $\varepsilon<\varepsilon_{AC}$, the ground state is a singlet state and we drive the transitions ${S}\rightarrow({T}_{-},{T}_{0},{T}_{+})$; when $\varepsilon>\varepsilon_{AC}$ the transitions are ${T}_{-}\rightarrow({S},{T}_{0},{T}_{+})$. When $\varepsilon<\varepsilon_{AC}$the Larmor frequency and the linewidth of the three transitions rapidly increases when reducing $\varepsilon$. When $\varepsilon>\varepsilon_{AC}$, the shifts in Larmor frequencies for the three transitions exhibit a diminishing trend due to the reduction of $J$ with $\varepsilon$. We also observe that the line-widths of all the transitions also become smaller with larger detuning (smaller exchange).
We next study the evolution of the Rabi frequencies as a function of detuning and magnetic field for each transition. Coherent Rabi oscillations between the ground state and the three excited states at $\varepsilon=2.3$ meV are measured and shown in Figs. S2 and S3. From the fast Fourier transform of those measurements, we extract the Rabi frequencies versus magnetic field, which are shown in Fig. 2(b). While the transitions to the ${T}_{0}$ and the ${T}_{+}$ states do not show any strong magnetic field dependence above 60 mT, this is strikingly different for the $T_{-}-S$ transition, which exhibits a non-monotonic behaviour with the highest Rabi frequency observed close to the position of the anticrossing, indicated by the dashed black line in Fig. 2(b). This behaviour contrasts with what is typically observed in systems with strong SOI, where the Rabi frequency scales linearly with the magnetic field  [35, 36, 37, 16]. By fixing the magnetic field to $B_{y}=90$ mT and plotting the Rabi frequency versus detuning, we observe that now all transitions show a decreasing Rabi frequency as the exchange interaction is reduced. This last observation indicates that exchange plays a vital role in the driving mechanism
@@
Figure 4: (a) Inhomogeneous dephasing time for the three transitions at in-plane field as a function of detuning. The colour code of each transition remains the same as previously described. Solid lines represent the fitted $T_{2}^{\ast}$ values for each transition, while the dashed lines indicate the dephasing time expected if only the detuning noise contribution is considered. The purple and green dashed lines overlap and the peak predicted by the red fit (corresponding to the anticrossing) is not observed experimentally. Each measurement for extracting $T_{2}^{\ast}$ has been integrated for 20 minutes. The fitting parameters can be found in the Supplementary. (b) Detuning susceptibility extracted from the derivative of the Larmor frequency of each transition with respect to $\varepsilon$. The orange colour refers to the out-of-plane $S-T_{0}$ transition at 12.5 mT. (c) Inhomogeneous dephasing time for the $S-T_{0}$ transition at out-of-plane field as a function of detuning, highlighting that the dephasing time is limited either by Zeeman or detuning noise. (d) Inhomogeneous dephasing time for the $S-T_{0}$ transition at $\varepsilon=7.4$ meV as a function of $B_{z}$, demonstrating an insensitivity to the magnetic field value. The error bars correspond to the standard deviation of the fit.
@@
In the in-plane regime, we obtain a larger $T_{2}^{\ast}$ and faster ac-driving due to a large SOI-induced tilt of quantisation axes of the two spins. This large value enables us to probe all possible transitions between different states, and produce the large exchange anisotropy of $\delta\theta=45^{\circ}$. Optimising $\delta g_{\perp}$ by controlling the quantum dots shapes and thus g-tensors, e.g. via machine learning algorithms  [42], will open up the path to leakage-protected two-qubit gates between ST-qubits [43], as well as to more efficient protocols for improving two-qubit gate fidelity between Loss-DiVincenzo qubits [32].
In the out-of-plane regime, we also demonstrate a much more uniform quantisation axis between the two dots and $\delta\theta\lesssim 5^{\circ}$, implying an almost isotropic exchange interaction. The properties of this regime enable a larger separation between the computational ($S,T_{0}$) and non-computational states ($T_{\pm}$) at low magnetic fields. This separation strongly suppresses leakage, allows uniform readout using conventional Pauli Spin Blockade, and maintains individual qubit addressability. Despite the use of natural germanium, which limits the dephasing time due to the larger effect of hyperfine noise in out-of-plane magnetic fields, we achieve a $T_{2}^{\ast}$ time of 400 ns. Importantly, this issue can be straightforwardly resolved in future devices by enrichment of isotopes with zero-spin, a method already common practice for silicon-based qubits. The coherence time is expected to increase one order of magnitude in enriched germanium and become comparable to state-of-the-art hole qubits in silicon [17]. These features along with the ac-driving demonstrated in this work will provide the key elements to scale up ST-based quantum processors [44, 45, 46].
@@ Methods
All measurements are performed in a Leiden cryogenics dry dilution cryostat with a base temperature of 50 mK. The fridge is equipped with a two-axis magnet along the Y and Z directions indicated in Fig. 1(a). A detailed sketch of the experimental setup is shown in Fig. S1
@@
We model the double quantum dot close to the transition $(1,1)\to(0,2)$ at detuning $\varepsilon\approx 0$ by a $5\times 5$ extended Fermi-Hubbard Hamiltonian spanning the states $(|S_{02}\rangle,|S_{11}\rangle,|T_{\uparrow}\rangle,|T_{\downarrow}\rangle,|T_{0}\rangle)$ [32, 47, 34]
@@
where $\underline{g}_{L/R}$ are the symmetric $g$ tensors of the left and right dots, $\underline{R}$ is a rotation matrix around the SOI axis by the spin-flip angle $\theta_{\text{so}}=2\arctan 2(t_{\text{sf}}/t_{\text{sc}})\approx 2d/l_{\text{so}}$, with $d$ being the distance between the dots and $l_{\text{so}}$ is the spin-orbit length. In the main text, the definitions of $\underline{g}_{1}$ and $\underline{g}_{2}$ refer to $\underline{g}_{L}\underline{R}(-\theta_{\text{so}}/2)$ and $\underline{g}_{R}\underline{R}(+\theta_{\text{so}}/2)$, respectively.
@@
We now introduce a more convenient singlet-triplet frame by first diagonalising the singlet sector by including the tunnelling $t$ via a rotation of an angle $\theta=-\text{arctan2}(\sqrt{2}t_{c}/\varepsilon)/2$. We also fix the direction of the spin quantization axis such that the triplet subsector is diagonal by performing a rotation $\underline{R}_{B}$ that maps $\bar{\textbf{b}}$ to the $z$-direction, i.e.
@@
These equations correspond to Eq. (1) and $H_{\text{driving}}$ in the main text, where the state $|S_{02}\rangle$ was not included. As the azimuthal angle $\varphi_{B}$ has no impact on the spectrum and amplitudes to the lowest order, there we set it to $\varphi_{B}=0$.
@@
When the magnetic field is large, it is convenient to change frame. Following  [32], we move to the qubit frame, defined by diagonalising the individual qubits. In this frame, we find the two-qubit Hamiltonian with anisotropic exchange interaction
@@
We now discuss a protocol to identify the relevant experimental parameters from the magnetic field sweep in Fig. 1(c) and 3(a). Here, we consider a fixed value of detuning $\varepsilon$ and of magnetic field direction, and vary only the amplitude of $B$.
@@
Rabi chevrons.
Here we show Rabi chevrons of the three studied transitions in the in-plane magnetic field direction.
Figure S2: (a-c) Rabi chevron for the 3 transitions at $B_{y}=40$ mT and $\varepsilon=4.0$ meV. (d-f) Fast Fourier transformation for the Rabi Chevrons.
@@
Figure S8: Additional measurements in the out-of-plane direction at 7 mT and $\varepsilon=4.0$ meV. a) Rabi Chevron showing microwave driven $S-T_{0}$ oscillations. b) Spectroscopy revealing the evolution of the three transistions with detuning. c) Zoom-in at low powers of the crossing between the S and $T_{-}$ states. The change in the background signal at 1.45 meV does not allow resolving the merging point of the two lines and therefore whether a crossing or an anticrossing takes place. However, if there is an anticrossing, it has a maximum size of 30 MHz.
@@
Here we present the measured Ramsey oscillations from which we extract the inhomogeneous dephasing time $T_{2}^{\ast}$ by fitting the data to the expression $\text{A}\text{cos}(\omega t)e^{-t/T_{2}^{\ast}}+\text{B}$. In this section we present the measurements and the fits from 4.

### v2

# Exchange anisotropies in microwave-driven singlet-triplet qubits

@@ Abstract
Hole spin qubits are rapidly emerging as the workhorse of semiconducting quantum processors because of their large spin-orbit interaction, enabling fast all-electric operations at low power. However, spin-orbit interaction also causes non-uniformities in devices, resulting in locally varying qubit energies and site-dependent anisotropies. While these anisotropies can be used to drive single-spins, if not properly harnessed, they can hinder the path toward large-scale quantum processors. Here, we report on microwave-driven singlet-triplet qubits in planar germanium and use them to investigate the anisotropy of two spins in a double quantum dot. We show two distinct operating regimes depending on the magnetic field direction. For in-plane fields, the two spins are largely anisotropic, and electrically tunable, which enables to measure all the available transitions; coherence times exceeding 3 $\mu$s are extracted. For out-of-plane fields, they have an isotropic response but preserve the substantial energy difference required to address the singlet-triplet qubit. Even in this field direction, where the qubit lifetime is strongly affected by nuclear spins, we find 400 ns coherence times. Our work adds a valuable tool to investigate and harness the anisotropy of spin qubits and can be implemented in any large-scale NxN device, facilitating the path towards scalable quantum processors.
@@
Here, we make use of a double quantum dot device formed in a Ge/SiGe heterostructure to characterize the exchange anisotropy, a key quantity for future scaling of semiconductor spin qubit quantum processors. We develop a simple protocol which allows to systematically characterize all the relevant spectroscopic features of two-hole systems at fixed magnetic field directions, including the angle between the quantization axes. For in-plane fields, we investigate microwave-driven singlet-triplet transitions –including the elusive $T_{-}$ to $T_{+}$ transition– and we show that such transitions are driven via the modulation of the exchange interaction and are just possible because of g-tensor dissimilarities. We further demonstrate that the quantization axes misalignment of the two spins can be electrically tuned. For out-of-plane fields, we demonstrate spin axes alignment and realize a $S-T_{0}$ microwave-driven qubit, which already for small fields has a large energy splitting and small transition matrix elements to the leakage states $T_{-}$ and $T_{+}$ [20]. This solves the problem of the non-orthogonal rotation axes, which have been limiting their implementation.
@@
We start by mapping out the amplitude of the signal at the Pauli Spin Blockade point as a function of the in-plane magnetic field $B_{y}$ and frequency of the microwave burst as seen in Fig. 1(c). We observe the appearance of three spin transitions, in contrast to the typically observed two lines  [26, 27, 28]. Those transitions are due to spin flips of charges localized in the two QDs. The faint line at the lowest frequency has half the frequency of the line indicated by the red arrow. This subharmonic transition is a consequence of a non-linear driving mechanism (see Fig. S11)  [29, 30].
@@
where $\theta=-\text{arctan2}(\varepsilon,\sqrt{2}t_{c})/2$ is the mixing angle, $J=-\varepsilon/2+\sqrt{\varepsilon^{2}/4+8t_{c}^{2}}$ is the exchange interaction, with $t_{c}$ the tunnel coupling. $\bm{b}$ ($\delta\bm{b}$) is the addition (difference) of the Zeeman vectors $\mu_{B}\underline{g}_{i}^{\text{so}}\bm{B}/2$ of the two dots ($i=1,2$), where $\mu_{B}$ is the Bohr magneton and $\bm{B}$ is the magnetic field. The parallel and perpendicular projections of $\delta\bm{b}$ on $\bm{b}$ are denoted $\delta b_{\parallel}$ and $\delta b_{\perp}$, respectively (see Fig. 1(f)). For convenience, we introduce the corresponding dimensionless g-factors as $\bar{b}=\mu_{B}|(\underline{g}_{1}^{\text{so}}+\underline{g}_{2}^{\text{so}})\bm{B}|/2=\mu_{B}\bar{g}|\bm{B}|$, $\delta b_{\parallel}=\mu_{B}\delta g_{\parallel}|\bm{B}|$ and $\delta b_{\perp}=\mu_{B}\delta g_{\perp}|\bm{B}|$. It is important to note that $\bar{g}$, $\delta g_{\parallel}$ and $\delta g_{\perp}$ are effective g-factors, which also incorporate the spin-flip tunnelling effects  [35, 33, 34].
@@
We emphasize that an equivalent description appropriate for large B field comprises two aligned spins coupled by an anisotropic exchange interaction matrix $JR_{y}(-\delta\theta)$. The rotation angle is given by $\delta\theta=\text{arctan}\left[\delta g_{\perp}/(\bar{g}+\delta g_{\parallel})\right]+\text{arctan}\left[\delta g_{\perp}/(\bar{g}-\delta g_{\parallel})\right]$ (see Methods). This angle is the misalignement between both quantization axes (see Fig. 1(f) and Fig. S2). From the above experimentally extracted values we find $\delta\theta=45^{\circ}$, highlighting the large degree of exchange anisotropy in our system. We note that large anisotropy is expected to be useful for resonant two-qubit gates in Loss-DiVincenzo qubits and for leakage protected two-qubit gates in ST qubits. Therefore, the ability to electrically tune this anisotropy could serve as a crucial control knob for qubit operation.
Figure 2: Electrical tunability of $\bar{g}$, $\delta g_{\perp}$, $\delta g_{\parallel}$ and the quantization axes tilt. (a, b) Magnetic field spectroscopy for two different voltage configurations of the electrostatic gates. The colored solid lines are the transitions corresponding to the parameters written in the legend. Panels (c, d) show the voltage differences of each configuration with respect to the voltage configuration 1 (Fig. 1 (c)). (c) Extracted parameters for each electrostatic configuration, including the misalignment $\delta\theta$, demonstrating electrical tunability.
Previous studies have independently demonstrated the electrical tunability of the g-tensor [36, 37], and the site-dependent properties between neighboring dots for a fixed electrostatic potential [11, 16]. To investigate the electrical tunability of the anisotropy of the dots, we tuned the DQD into two additional voltage configurations and repeated the measurement shown Fig. 1(c). Figure 2 (a, b) shows the magnetic field spectroscopy at those two different electrostatic potentials; the $T_{-}-S$, $T_{-}-T_{0}$ and $T_{-}-T_{+}$ are again observed. As a consequence of the change in the electrostatic configuration (Figs. 2 (c, d)), the slopes of the transitions are different, implying that the tilt between the quantization axes changes. In voltage configuration 2 (Fig. 2 (a)), the transitions $T_{-}-S$ and $T_{-}-T_{0}$ are almost parallel, indicating that the Zeeman energy difference $\delta b_{\parallel}$ is small and that the $S-T_{0}$ splitting is dominated by exchange. On the other hand, the slopes in voltage configuration 3, (Fig. 2 (b)) are smaller and dissimilar. Figure 2(e) compares the extracted parameters for the different voltage configurations. By calculating the relative tilt, we show the tunability of the misalignment by 10 degrees with gate voltage. This finding illustrates not only the utility of this protocol in quickly characterizing the site-dependent anisotropies, without extracting each g-tensor, but also the ability to control these anisotropies electrically.
@@
After the characterization of the spectroscopic features, we return to the voltage configuration 1 and focus on extracting the mechanism which permits to drive these transitions. We study the dependence of the spin transitions as a function of detuning at a fixed magnetic field in Fig. 3(a). As the system is initialized in the ground state, the three observed lines correspond to the transitions indicated by the three arrows depicted in Fig. 1(e). From the $T_{-}-S$ transition, we can experimentally determine the position, $\varepsilon_{AC}=1.8$ meV, and size, $\Delta_{ST_{-}}=175$ MHz, of the $S-T_{-}$ anticrossing. As this point sets the turnover of the ground state, we identify two regimes. When $\varepsilon<\varepsilon_{AC}$, the ground state is a singlet state and we drive the transitions ${S}\rightarrow({T}_{-},{T}_{0},{T}_{+})$; when $\varepsilon>\varepsilon_{AC}$ the transitions are ${T}_{-}\rightarrow({S},{T}_{0},{T}_{+})$. When $\varepsilon<\varepsilon_{AC}$, the Larmor frequency and the linewidth of the three transitions rapidly increases when reducing $\varepsilon$. When $\varepsilon>\varepsilon_{AC}$, the shifts in Larmor frequencies for the three transitions exhibit a diminishing trend due to the reduction of $J$ with $\varepsilon$. We also observe that the line-widths of all the transitions also become smaller with larger detuning (smaller exchange).
We next study the evolution of the Rabi frequencies as a function of detuning and magnetic field for each transition. Coherent Rabi oscillations between the ground state and the three excited states at $\varepsilon=2.3$ meV are measured and shown in Figs. S3 and S5 allowing us to estimate the gate fidelities (see S4). From the fast Fourier transform of those measurements, we extract the Rabi frequencies versus magnetic field, which are shown in Fig. 3(b). While the transitions to the ${T}_{0}$ and the ${T}_{+}$ states do not show any strong magnetic field dependence above 60 mT, this is strikingly different for the $T_{-}-S$ transition, which exhibits a non-monotonic behaviour with the highest Rabi frequency observed close to the position of the anticrossing, indicated by the dashed black line in Fig. 3(b). This behaviour contrasts with what is typically observed in systems with strong SOI, where the Rabi frequency scales linearly with the magnetic field  [38, 39, 36, 16]. By fixing the magnetic field to $B_{y}=90$ mT and plotting the Rabi frequency versus detuning, we observe that now all transitions show a decreasing Rabi frequency as the exchange interaction is reduced. This last observation indicates that exchange plays a vital role in the driving mechanism.
@@
Figure 5: (a) Inhomogeneous dephasing time for the three transitions at in-plane field as a function of detuning. The colour code of each transition remains the same as previously described. Solid lines represent the fitted $T_{2}^{\ast}$ values for each transition, while the dashed lines indicate the dephasing time expected if only the detuning noise contribution is considered. The purple and green dashed lines overlap and the peak predicted by the red fit (corresponding to the anticrossing) is not observed experimentally. Each measurement for extracting $T_{2}^{\ast}$ has been integrated for 20 minutes. The fitting parameters can be found in the Supplementary. (b) Detuning susceptibility extracted from the derivative of the Larmor frequency of each transition with respect to $\varepsilon$. The orange colour refers to the out-of-plane $S-T_{0}$ transition at 12.5 mT. The inset shows a zoom-in at large detunings. (c) Inhomogeneous dephasing time for the $S-T_{0}$ transition at out-of-plane field as a function of detuning, highlighting that the dephasing time is limited either by Zeeman or detuning noise. (d) Inhomogeneous dephasing time for the $S-T_{0}$ transition at $\varepsilon=7.4$ meV as a function of $B_{z}$, demonstrating an insensitivity to the magnetic field value. The error bars correspond to the standard deviation of the fit.
@@
In the in-plane regime, we obtain a larger $T_{2}^{\ast}$ and faster ac-driving due to a large SOI-induced tilt of quantization axes of the two spins. This large value enables us to probe all possible transitions between different states, and produce the large exchange anisotropy of $\delta\theta=45^{\circ}$. In addition, we show that $\delta\theta$ is electrically tunable and therefore optimizing $\delta g_{\perp}$ by controlling the quantum dots shapes and thus g-tensors, e.g. via machine learning algorithms  [44], will open up the path to leakage-protected two-qubit gates between ST-qubits [45, 46, 4], as well as to more efficient protocols for improving two-qubit gate fidelity between Loss-DiVincenzo qubits [33].
In the out-of-plane regime, we also demonstrate a much more uniform quantization axis between the two dots and $\delta\theta\lesssim 5^{\circ}$, implying an almost isotropic exchange interaction. The properties of this regime enable a larger separation between the computational ($S,T_{0}$) and non-computational states ($T_{\pm}$) at low magnetic fields. This separation strongly suppresses leakage, allows uniform readout using conventional Pauli Spin Blockade, and maintains individual qubit addressability. Despite the use of natural germanium, which limits the dephasing time due to the larger effect of hyperfine noise in out-of-plane magnetic fields, we achieve a $T_{2}^{\ast}$ time of 400 ns. Importantly, this issue can be straightforwardly resolved in future devices by enrichment of isotopes with zero-spin, a method already common practice for silicon-based qubits. The coherence time is expected to increase one order of magnitude in enriched germanium and become comparable to state-of-the-art hole qubits in silicon [17]. These features along with the ac-driving demonstrated in this work will provide the key elements to scale up ST-based quantum processors [46, 47, 48].
@@ Methods
All measurements are performed in a Leiden cryogenics dry dilution cryostat with a base temperature of 50 mK. The fridge is equipped with a two-axis magnet along the Y and Z directions indicated in Fig. 1(a). A detailed sketch of the experimental setup is shown in Fig. S1.
@@
We model the double quantum dot close to the transition $(1,1)\to(0,2)$ at detuning $\varepsilon\approx 0$ by a $5\times 5$ extended Fermi-Hubbard Hamiltonian spanning the states $(|S_{02}\rangle,|S_{11}\rangle,|T_{+}\rangle,|T_{-}\rangle$, $|T_{0}\rangle)$ [33, 49, 35]
@@
where $\underline{g}_{L/R}$ are the symmetric $g$ tensors of the left and right dots, $\underline{R}$ is a rotation matrix around the SOI axis by the spin-flip angle $\theta_{\text{so}}=2\arctan 2(t_{\text{sf}},t_{\text{sc}})\approx 2d/l_{\text{so}}$, with $d$ being the distance between the dots and $l_{\text{so}}$ is the spin-orbit length. In the main text, the definitions of $\underline{g}_{1}^{\text{so}}$ and $\underline{g}_{2}^{\text{so}}$ refer to $\underline{g}_{L}\underline{R}(-\theta_{\text{so}}/2)$ and $\underline{g}_{R}\underline{R}(+\theta_{\text{so}}/2)$, respectively (see Fig. S2(a)).
@@
We now introduce a more convenient singlet-triplet frame by first diagonalising the singlet sector by including the tunnelling $t$ via a rotation of an angle $\theta=-\text{arctan2}(\varepsilon,\sqrt{2}t_{c})/2$. We also fix the direction of the spin quantization axis such that the triplet subsector is diagonal by performing a rotation $\underline{R}_{B}$ that maps $\bar{\textbf{b}}$ to the $z$-direction, i.e.
@@
These equations correspond to Eq. (1) and $H_{\text{driving}}$ in the main text, where the state $|S_{02}\rangle$ was not included as its energy contribution can be neglected. As the azimuthal angle $\varphi_{B}$ has no impact on the spectrum and amplitudes to the lowest order, there we set it to $\varphi_{B}=0$.
@@
We note the second order coupling terms $\delta\varepsilon\delta b_{\perp}$ and $\delta\varepsilon\delta b_{\parallel}$ drive the Rabi oscillations despite the diagonal structure of the triplet sector in the Hamiltonian.
@@
When the magnetic field is large, it is convenient to change frame. Following  [33], we move to the qubit frame, defined by diagonalizing the individual qubits. In this frame, we find the two-qubit Hamiltonian with anisotropic exchange interaction
@@
We now discuss a protocol to identify the relevant experimental parameters from the magnetic field sweep in Fig. 1(c) and 4(a). Here, we consider a fixed value of detuning $\varepsilon$ and of magnetic field direction, and vary only the amplitude of $B$.
@@
Geometrical interpretation.
Figure S2: (a) Schematic explaining the difference between the laboratory and the spin-orbit frames. The light grey/brown arrows correspond to the g-tensors of left/right quantum dots, respectively. The dark grey/brown arrows represent the quantization axes in the spin-orbit frame, where the light quantization axes have been rotated by an angle $\theta_{\text{so}}/2$ in order to take into account the spin-flip tunneling. The angle $\delta\theta$ between quantization axes 1 and 2 considers the spin-flip tunneling and the tilt between the g-tensors $\underline{g}_{L}$ and $\underline{g}_{R}$. (b) Geometrical explanation of $\delta\theta$ using the parameters extracted from the spectroscopy as a function of the magnetic field. By applying the trigonometric relations one can prove that $\delta\theta=\text{arctan}\left[\delta g_{\perp}/(\bar{g}+\delta g_{\parallel})\right]+\text{arctan}\left[\delta g_{\perp}/(\bar{g}-\delta g_{\parallel})\right]$.
Rabi chevrons and gate fidelities.
Here we show Rabi chevrons of the three studied transitions in the in-plane magnetic field direction and the estimation of the gate fidelities.
Figure S3: (a-c) Rabi chevron for the 3 transitions at $B_{y}=40$ mT and $\varepsilon=4.0$ meV. (d-f) Fast Fourier transformation for the Rabi Chevrons. The distinction between the $T_{-}$ and $T_{+}$ states is made possible by the Landau-Zener transition probability when returning to the readout point. In our system, the $T_{-}$ state in the effective (1,1) configuration maps into an unblocked state due to the anticrossing, while the other states remain blocked.
Figure S4: Rabi oscillations for the three transitions investigated at different magnetic field and detuning point. Each trace is fitted to: $A\cos(2\pi f_{R}t+\phi)\exp(-t/T_{2}^{\text{Rabi}})+B$. Data (a-c) are the resonant traces from the Rabi chevrons in Fig.S3. (d-f) are the line traces at 150 mT from the data shown in Fig. S5.The gate fidelity of each transition is estimated using $F=1/2(1+exp(-T_{2}^{\text{Rabi}}f_{R}/2))$, where $T_{2}^{\text{Rabi}}$ the Rabi decay and $f_{R}$ the Rabi frequency [50].
@@
Figure S10: Additional measurements in the out-of-plane direction at 7 mT and $\varepsilon=4.0$ meV. a) Rabi Chevron showing microwave driven $S-T_{0}$ oscillations. b) Spectroscopy revealing the evolution of the three transistions with detuning. c) Zoom-in at low powers of the crossing between the S and $T_{-}$ states. The change in the background signal at 1.45 meV does not allow resolving the merging point of the two lines and therefore whether a crossing or an anticrossing takes place, but it allows to set the upper bound for the anticrossing size at 30 MHz. The fact that there is a blocked return probability at 1.45 meV demonstrates that there is an anticrossing where the $S$ and $T_{-}$ states mix.
@@
Here we present the measured Ramsey oscillations from which we extract the inhomogeneous dephasing time $T_{2}^{\ast}$ by fitting the data to the expression $\text{A}\text{cos}(\omega t)e^{-t/T_{2}^{\ast}}+\text{B}$. In this section we present the measurements and the fits from 5.